OpenStudy (anonymous):

what is the solution to the system of linear equations below? {x + y = 5 {x - y = 3

jimthompson5910 (jim_thompson5910):

One way to solve is to add the equations (straight down) x + y = 5 x - y = 3 ---------- 2x+0y = 8 Notice how the two x's add to 2x, the y and -y add to 0y which goes away, and the 5 & 3 on the right side combine to 5+3 = 8 So we end up with 2x = 8 which means x = ???

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

these are the answer choices a (8,-3) b (6,-1) c (5,2) d (4,1)

jimthompson5910 (jim_thompson5910):

if 2 times _____ = 8, then what goes in the blank?

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

@jim_thompson5910